OpenStudy (anonymous):

can someone help me with this problem? "use the change of base formula to evaluate each expression. Then convert it to a logarithm in base 8": log(base2)9

OpenStudy (anonymous):

i got \[\log_{8} 729\] is that correct?

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

use the formula log(base b)x=log(base a)x/log(base a)b log(base2)9=log(base8)9/log(base8)2, now you can still simplify this by dividing them

OpenStudy (anonymous):

Let \(k = log_2{9} \implies 2^k = 9\) \[ \implies log_8 (2^k) = log_8 9\] \[ \implies k\ log_8 2 = log_8 9\] \[ \implies k = \frac{log_8 9}{log_8 2}\] \[ \implies log_2 9 = \frac{log_8 9}{log_8 2}\]